OpenStudy (anonymous):

What is the 24th term in -5,-15,-45,-135,-405...

Parth (parthkohli):

Sure I'd help :)

Parth (parthkohli):

\( \color{Black}{\Rightarrow -5 \cdot 3^{22} }\) Seems like we'd have a huge number

OpenStudy (anonymous):

Thanks I cant remember the equation for it and yeah its suppose to be a big number

Parth (parthkohli):

Oops it's 3 * 3^(23)

OpenStudy (anonymous):

haha Thanks!

Parth (parthkohli):

Yeah, the nth term in a geometric sequence is \(a \times r^{n - 1}\)

OpenStudy (anonymous):

yeah but how should you solve it using that?

OpenStudy (anonymous):

i cant plug the numbers in

Parth (parthkohli):

a is the first term and r is the common ratio. \( \color{Black}{\Rightarrow -5 \times 3^{24 - 1}}\)

Parth (parthkohli):

n is the nth term

OpenStudy (anonymous):

the what term?

Parth (parthkohli):

We want the 24th term, so we'll plug in n with 24

Parth (parthkohli):

Got it?

OpenStudy (anonymous):

yeah but y page is showing me the answers could be 94,143,178,827 -470715894135

OpenStudy (anonymous):

or

OpenStudy (anonymous):

-1412147682405 -41943040

Parth (parthkohli):

Oh lol google

OpenStudy (anonymous):

i tried

Parth (parthkohli):

It has the answer on top

OpenStudy (anonymous):

oh okay! Great Thanks!

Parth (parthkohli):

You're welcome!

Parth (parthkohli):

If you think my answer was good, you can click on the "Best Answer" button :)